OpenStudy (anonymous):

Solve the given linear system 2x-4y=-6 -x+2y=3

OpenStudy (anonymous):

Just sum the equations !:)

OpenStudy (anonymous):

you can solve by elimination so you multiply the 2nd equation by 2 to get -2x+4y=6 so you get 2x-4y=-6 -2x+4y=6 0=0 since both sides are the same the answer is all solutions meaning they are the same line

OpenStudy (anonymous):

But after that find a way to kill one of the X s or Y s ! @insanitycat said it for you with kill X s complate !

OpenStudy (anonymous):

uhmm.. could you provide example @E.ali

OpenStudy (anonymous):

Sure ! Look care fully !

OpenStudy (anonymous):

sure..

OpenStudy (anonymous):

If we have : 9x-7y=1 5x+7y=3 there Y s are Symmetric . So if we sum the equations have : 9x+5x+7y-7y=1+3 have : 14x+0=4 14x=4 x=4/14 Got it ?:)

OpenStudy (texaschic101):

this can also be solved with substitution

OpenStudy (anonymous):

@texaschic101 :Yes ! Thank you !

OpenStudy (texaschic101):

-x + 2y = 3 -x = -2y + 3 (multiply by -1 to make x positive) x = 2y - 3 now sub 2y - 3 in for x in the other equation and solve for y. Once y is found, sub that back into either equation and solve for x.

OpenStudy (anonymous):

but @texaschic101 you see. 2x-4y=-6 2(2y-3)-4y=-6 4y-6-4y=-6 -6=-6

OpenStudy (anonymous):

so y=0?

OpenStudy (texaschic101):

no.... 2x - 4y = -6 2(2y - 3) - 4y = -6 4y - 6 - 4y = -6 (add 6 to both sides) 4y - 4y = -6 + 6 0 = 0 if you get a problem where the end result is EQUAL, the the solution to this problem is INFINITE SOLUTIONS now if the problem would have ended 0 = 3 or 2 = 4, these are incorrect and the answer for this would have been no solution.

OpenStudy (anonymous):

so y=0?

OpenStudy (texaschic101):

The answer to this problem is INFINITE SOLUTIONS.

OpenStudy (anonymous):

ooh.. i see.. thanks :))

OpenStudy (texaschic101):

its ok :)

OpenStudy (anonymous):

thank you very much :"> btw, how young are you?